# RF4D:Neural Radar Fields for Novel View Synthesis in Outdoor Dynamic Scenes

### v2

@@ 4.3 Training objective
Our training objective combines three losses: (1) a radar power reconstruction loss supervising the rendered measurement, (2) a temporal consistency regularization over occupancy, and (3) a sparsity-inducing penalty that prevents trivial solutions.
In each iteration, we sample $N$ range-azimuth bins and each range-azimuth bin $(\delta,\theta)$ receives predicted occupancy $\hat{\alpha}_{\delta,\theta}$ and RCS $\hat{\sigma}_{\delta,\theta}$, and the rendered power is $\hat{P}_{\delta,\theta}=\hat{\alpha}_{\delta,\theta}\cdot\log_{10}(\hat{\sigma}_{\delta,\theta}/\delta^{2})$.
@@
To encourage temporal coherence, we ensure that the predicted occupancy field remains consistent across adjacent time frames. Let $\hat{\alpha}^{t-\Delta t}_{\delta,\theta}$ and $\hat{\alpha}^{t+\Delta t}_{\delta,\theta}$ be the predicted occupancies of the same bin at neighboring frames. The temporal consistency loss is then:
$\mathcal{L}_{\text{oc}}=\frac{1}{N}\sum_{\delta,\theta}\left(\left(\hat{\alpha}_{\delta,\theta}-\hat{\alpha}^{t-\Delta t}_{\delta,\theta}\right)^{2}+\left(\hat{\alpha}_{\delta,\theta}-\hat{\alpha}^{t+\Delta t}_{\delta,\theta}\right)^{2}\right).$ (8)
@@
The final objective combines these three terms:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rt}}+\lambda_{\text{oc}}\mathcal{L}_{\text{oc}}+\lambda_{\text{p}}\mathcal{L}_{\text{p}},$ (10)
where $\lambda_{\text{oc}}$ and $\lambda_{\text{p}}$ control the contributions of temporal regularization and sparsity.
@@ 5 Experiments
Table 1: Quantitative comparison across three different driving scenarios from the RobotCar dataset [50]. The best result and the runner-up are highlighted in bold and underline, respectively.
Scene 1 Scene 2 Scene 3
Method PSNR↑ RMSE↓ CD↓ RCD↓ PSNR↑ RMSE↓ CD↓ RCD↓ PSNR↑ RMSE↓ CD↓ RCD↓
D-NeRF [41] 20.98 0.0894 39.6776 0.0862 22.31 0.0861 28.6662 0.1229 20.51 0.0945 139.5111 0.3127
DyNeRF [46] 20.74 0.0919 59.9361 0.0776 20.60 0.0934 30.1186 0.0743 20.55 0.0941 66.7763 0.5162
RadarFields [20] 21.15 0.0885 16.0938 0.0176 20.99 0.0901 10.4556 0.0107 14.83 0.2318 7.2165 0.0286
Ours 22.95 0.0715 3.6765 0.0083 23.06 0.0704 2.7455 0.0110 23.25 0.0697 4.2011 0.0109
Table 2: Quantitative comparison across three driving scenarios from the Boreas dataset [51]. The best result and the runner-up are highlighted in bold and underline, respectively.
Scene 1 Scene 2 Scene 3
Method PSNR↑ RMSE↓ CD↓ RCD↓ PSNR↑ RMSE↓ CD↓ RCD↓ PSNR↑ RMSE↓ CD↓ RCD↓
D-NeRF [41] 22.87 0.0719 50.8978 0.0830 22.48 0.0752 46.1929 0.1852 21.07 0.0885 204.4935 0.6521
DyNeRF [46] 22.10 0.0786 13.9267 0.0547 22.33 0.0766 93.9153 0.2306 21.30 0.0862 19.5226 0.0353
RadarFields [20] 13.76 0.2080 17.1700 0.1003 23.13 0.0708 13.6600 0.0273 19.99 0.1036 9.0822 0.0142
Ours 23.94 0.0637 13.2808 0.0244 24.21 0.0617 12.7193 0.0166 23.45 0.0677 7.4702 0.0130
Scene 1 Scene 2 Scene 3
Time Flow PSNR ↑ RMSE ↓ CD↓ RCD↓ PSNR ↑ RMSE↓ CD↓ RCD↓ PSNR↑ RMSE↓ CD↓ RCD↓
22.51 0.0752 5.9005 0.0200 22.31 0.0767 6.7077 0.0170 21.72 0.0831 7.1924 0.0114
22.92 0.0717 4.4440 0.0102 23.03 0.0707 2.7964 0.0102 22.98 0.0718 6.1560 0.0084
22.95 0.0715 3.6765 0.0083 23.06 0.0704 2.7455 0.0110 23.25 0.0697 4.2011 0.0109
Table 4: Ablation study for the Boreas [51] dataset. The best result and the runner-up are highlighted in bold and underline, respectively.
Scene 1 Scene 2 Scene 3
Time Flow PSNR ↑ RMSE ↓ CD↓ RCD↓ PSNR ↑ RMSE↓ CD↓ RCD↓ PSNR↑ RMSE↓ CD↓ RCD↓
22.49 0.0743 18.7007 0.0254 23.88 0.0640 18.1834 0.0167 23.05 0.0707 8.5524 0.0128
23.59 0.0663 15.1568 0.0283 24.03 0.0630 12.9487 0.0168 23.25 0.0692 7.3158 0.0129
23.94 0.0637 13.2808 0.0244 24.21 0.0617 12.7193 0.0166 23.45 0.0677 7.4702 0.0130
@@ 5.1 Experimental setup
We evaluate RF4D on two publicly available radar datasets: Oxford Radar RobotCar [50] and Boreas [51]. Both datasets contain spinning FMCW radar measurements capturing full 360-degree range-azimuth scans, synchronized with auxiliary sensors including GPS/IMU, cameras, and LiDAR. Accurate vehicle odometry is provided, enabling precise radar pose estimation for each frame. The datasets feature multiple driving scenarios with moving vehicles, making them particularly suitable for benchmarking radar-based novel view synthesis in outdoor dynamic scenes. For each dataset, we randomly select three scenes and each scene contains approximately 60–100 consecutive radar scans, from which 80% are randomly assigned as training data, with the remaining 20% held out as the test set.
@@
Following the preprocessing strategy outlined in RadarFields [20], we discard the innermost 69 range bins (approximately a 3-meter radius around the radar sensor) from all radar measurements. These bins predominantly reflect the metallic rooftop of the data collection vehicle and thus do not contribute meaningful scene information. We retain the radar measurements up to the maximum range bin of 1200 (approximately 50 meters), resulting in 1132 effective range bins. All azimuth angles (totaling 400) are retained without modification, yielding a final size of $(400,1132)$ per radar measurement. Additionally, each scene’s coordinates are uniformly scaled to fit within a normalized cubic volume $[-1,1]^{3}$, and timestamps are similarly normalized to the range $[-1,1]$.
@@
Our RF4D model is implemented in PyTorch [52] and trained on a single NVIDIA RTX A5000 GPU. For efficient neural field optimization, we use the tiny-cuda-nn framework [53] to implement neural network backbones and input encodings. The model is optimized using the Adam optimizer [54] with an initial learning rate of $5\times 10^{-4}$, which is progressively decayed by a factor of 0.1 during training, resulting in a final learning rate of $5\times 10^{-5}$. Each scene is trained independently for 6000 epochs. In each training iteration, we randomly sample a batch of 10 radar scans from the input sequence and further randomly select range-azimuth bins from these frames for supervision.
@@
In line with prior work [20], we evaluate the quality of novel radar measurement synthesis using Peak Signal-to-Noise Ratio (PSNR) and Root Mean Square Error (RMSE), computed between predicted and ground truth radar measurements. To assess occupancy estimation accuracy, we compute Chamfer Distance (CD) and Relative Chamfer Distance (RCD) between the predicted 2D Bird’s-Eye View (BEV) point clouds and ground truth point clouds. Since accurate ground-truth occupancy is difficult to extract directly from radar due to low azimuth resolution, multipath interference, and inherent sensor noise, we leverage synchronized LiDAR point clouds to construct geometric occupancy references. For each radar frame, we calibrate the corresponding LiDAR point clouds to radar coordinates using the provided LiDAR-to-radar transformation matrix. We then select points within the radar’s sensing range bounds and with $-1<z<1$. The resulting points are projected onto the $xy$-plane to generate the BEV ground-truth occupancy for that frame.
Note that to ensure compatibility with dynamic scenes ,we perform occupancy evaluation using only the LiDAR points clouds corresponding to each radar timestamp, without accumulating across time.
@@ 5.2 Comparison with state-of-the-art
To ensure fair and rigorous evaluation, we focus exclusively on NF-based approaches, excluding methods relying on handcrafted environment modeling or sensor-specific signal simulations.We compare RF4D with three representative state-of-the-art methods: Radar Fields [20], D-NeRF [41], and DyNeRF [46]. For Radar Fields, we directly adopt their official implementation, ensuring consistency and fair comparison. For D-NeRF and DyNeRF, originally designed for RGB-based dynamic NVS, we adapt their architectures to our radar-based pipeline and retrain accordingly.
@@
Tables 1 and 2 present quantitative comparisons on the Oxford Radar RobotCar and Boreas datasets, respectively. RF4D consistently achieves the best performance across all driving scenarios on both datasets, outperforming all competing methods in both radar measurement synthesis (PSNR, RMSE) and occupancy estimation (CD, RCD). Notably, on the Oxford dataset, our method achieves substantial improvements in occupancy accuracy, with CD and RCD reduced by large margins compared to competing approaches. Methods using volume rendering such as D-NeRF and DyNeRF result poor synthesis quality and occupancy accuracy. This suggests that such methods, originally designed for RGB inputs, are ill-suited for radar data due to their lack of radar-specific physical modeling. Among methods that incoporate radar sensing physics, our approach further outperforms Radar Fields across all scenes. We attribute this to RF4D’s incorporation of temporal information and temporal regularization, enabling it to better model dynamic scenes and produce more coherent predictions over time.
Figure 3: Qualitative comparison of novel-view radar measurement synthesis (top) and occupancy estimation (bottom) across different methods on the RobotCar dataset [50] . Ground truth occupancy is derived from LiDAR point clouds. Our method accurately reconstructs the radar measurements with clear structure and minimal artifacts, while RadarFields introduces noise and blurring, especially around dynamic objects. D-NeRF and DyNeRF, which rely on volume rendering, fail to recover meaningful scene structure. In contrast, our predictions are clean and closely matching the ground truth.
@@
Figure 3 presents qualitative comparisons (two different scenes) of reconstructed radar measurements and occupancy maps using different methods. Our approach synthesizes high-fidelity novel views with fine structural detail and accurate reflections, while competing methods struggle significantly. Methods using volume rendering, such as D-NeRF and DyNeRF, fail to recover meaningful scene structures due to a fundamental mismatch between their optical rendering assumptions and the physics of radar sensing. For occupancy estimation, D-NeRF and DyNeRF produce coarse estimations, missing critical details like object contours and road boundaries. Among radar-specific methods, our model outperforms Radar Fields significantly. Our reconstructed radar measurements exhibit fewer artifacts and our estimated occupancy maps closely match LiDAR ground truth, capturing both static structures and moving objects with high spatial fidelity. Radar signals are often interfered by multipath effects and internal sensor noise, leading to false-positive responses. Radar Fields tends to misinterpret these artifacts as occupied regions, leading to overestimation. In contrast, we employ a Gumbel-Sigmoid activation function for occupancy prediction, enforcing binary outputs and effectively suppressing uncertain responses. Moreover, our method maintains precise localization of moving objects, such as on-street vehicles (dots between the street lines in the ground truth occupancy map). Whereas Radar Fields produces blurred predictions for dynamic targets. This improvement is due to our inclusion of temporal information and the feature-level flow module to ensure temporal coherence. These results show the effectiveness of our radar-centric design, which integrates physically grounded signal modeling and temporal alignment, yielding robust performance in both novel-view measurement synthesis and occupancy estimation.
@@ 5.3 Ablation study
In Table 3 and Table 4, we further conduct ablation studies to validate the effectiveness of incorporating the time input and the feature-level flow module for temporal regularization. All methods listed here utilize our proposed radar-specific power rendering. When time is additionally included as input (second row), we further observe consistent improvements across all metrics and scenes. This indicates that temporal input is essential for accurate novel radar measurement synthesis under dynamic driving environments. Further introducing the feature-level flow module (third row), which learns residual offsets in latent space to align temporal features, yields more improvements (CD and RCD) in occupancy estimation. Additionally, PSNR and RMSE improve, indicating that the quality of reconstructed radar measurements also benefits from more accurate occupancy estimation. Overall, these results validate that explicitly incorporating time input and the feature-level flow module are both critical. Their combination achieves superior reconstruction quality and occupancy accuracy, justifying each component of our RF4D design.

### v3

@@ 4.3 Training objective
The training objective combines four components: a radar power reconstruction loss, a temporal consistency regularization, an occupancy sparsity regularization, and a motion offset regularization. During each iteration, $N$ range-azimuth bins $(\delta,\theta)$ are sampled. Each range-azimuth bin receives predicted occupancy $\hat{\alpha}_{\delta,\theta}$, RCS $\hat{\sigma}_{\delta,\theta}$, and motion offsets $\Delta x_{\delta,\theta}^{-}$ and $\Delta x_{\delta,\theta}^{+}$. The rendered power is $\hat{P}_{\delta,\theta}=\hat{\alpha}_{\delta,\theta}\cdot\log_{10}(\hat{\sigma}_{\delta,\theta}/\delta^{2})$.
@@
To encourage temporal coherence, the predicted occupancy at time $t$ is encouraged to remain consistent with its counterparts in the adjacent frames. This is achieved through a temporal regularization term:
$\mathcal{L}_{\text{oc}}=\frac{1}{N}\sum_{\delta,\theta}\left(\left(\hat{\alpha}_{\delta,\theta}-\hat{\alpha}^{t-\Delta t}_{\delta,\theta}\right)^{2}+\left(\hat{\alpha}_{\delta,\theta}-\hat{\alpha}^{t+\Delta t}_{\delta,\theta}\right)^{2}\right),$ (8)
where $\hat{\alpha}^{t-\Delta t}_{\delta,\theta}$ and $\hat{\alpha}^{t+\Delta t}_{\delta,\theta}$ are the predicted occupancies of the same point at adjacent frames obtained through Equation 5.
@@
Finally, the predicted motion offsets are regularized to prevent unrealistic displacements. We apply a regularization on the magnitude of offsets:
[…]
The final objective combines all terms:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rt}}+\lambda_{\text{oc}}\mathcal{L}_{\text{oc}}+\lambda_{\text{p}}\mathcal{L}_{\text{p}}+\lambda_{\text{m}}\mathcal{L}_{\text{m}},$ (11)
where $\lambda{\text{oc}}$, $\lambda_{\text{p}}$, and $\lambda_{\text{m}}$ are weighting factors that balance temporal regularization, sparsity, and motion constraints.
Figure 4: Qualitative comparison of novel-view radar measurement synthesis and occupancy estimation on the Oxford Radar RobotCar dataset [2]. Ground-truth occupancy is derived from LiDAR point clouds. RF4D reconstructs radar measurements with clear structures and preserved dynamic targets (red boxes), while Radar Fields produces noisier and blurrier results.
Table 5: Comparison of occupancy estimation results across different driving scenarios on the Oxford Radar RobotCar [2] and Boreas [6] datasets. The best and second-best results are highlighted in bold and underline, respectively.
Oxford RobotCar Boreas
Scene 1 Scene 2 Scene 3 Scene 4 Sun Rain Static
Method CD↓ RCD↓ CD↓ RCD↓ CD↓ RCD↓ CD↓ RCD↓ CD↓ RCD↓ CD↓ RCD↓ CD↓ RCD↓
CFAR [45] 19.5299 0.0276 30.5561 0.0387 7.9197 0.0185 26.4644 0.0291 8.7570 0.0214 8.7636 0.0234 13.7390 0.0083
Bayesian filtering [62] 13.2786 0.0179 21.7657 0.0273 6.3696 0.0208 18.6878 0.0206 6.0716 0.0150 8.6918 0.0069 8.4866 0.0049
Ours 7.3449 0.0190 8.0909 0.0125 3.2896 0.0097 18.7607 0.0147 5.4482 0.0033 10.4558 0.0051 8.0739 0.0040
@@ 5.1 Experimental setup
We evaluate RF4D on two publicly available radar datasets: Oxford Radar RobotCar [2] and Boreas [6]. Both datasets contain spinning FMCW radar measurements capturing full 360-degree range-azimuth scans, synchronized with auxiliary sensors including GPS/IMU, cameras, and LiDAR. Accurate vehicle odometry is provided, enabling precise radar pose estimation for each frame. The datasets cover diverse driving scenarios with moving vehicles, making them suitable for evaluating radar-based novel view synthesis in dynamic outdoor environments. For the Boreas dataset, we select four representative scenes: sun, snow, rain, and static. The sun, snow, and rain scenes contain dynamic traffic, while static has no moving objects. The snow and rain scenes are used to assess performance under adverse weather conditions. For the Oxford Radar RobotCar dataset, which only includes clear-weather data, we randomly choose four scenes. Each scene consists of approximately 60–100 consecutive radar scans, among which four scans are uniformly sampled for testing, and the remaining scans are used for training.
@@
Following the preprocessing strategy outlined in RadarFields [4], we discard the innermost 50 range bins (approximately a 3-meter radius around the radar sensor) from all radar measurements. These bins predominantly reflect the metallic rooftop of the data collection vehicle and thus do not contribute meaningful scene information. We retain the radar measurements up to the maximum range bin of 1200 (approximately 70 meters), resulting in 1150 effective range bins. All azimuth angles (totaling 400) are retained without modification, yielding a final size of $(400,1150)$ per radar measurement. Additionally, each scene’s coordinates are uniformly scaled to fit within a normalized cubic volume $1,1]^{3}$, and timestamps are similarly normalized to the range $[0,1]$.
@@
Our RF4D model is implemented in PyTorch [42] and trained on a single NVIDIA RTX A5000 GPU. For efficient neural field optimization, we use the tiny-cuda-nn framework [38] to implement neural network backbones and input encodings. The model is optimized using the Adam optimizer [28] with an initial learning rate of $1\times 10^{-4}$, which is progressively decayed by a factor of 0.1 during training, resulting in a final learning rate of $1\times 10^{-5}$. Each scene is trained independently for 15000 iterations. In each training iteration, we randomly sample a batch of 4 radar scans from the input sequence and further randomly select range-azimuth bins from these frames for supervision.
@@
In line with prior work [4], we evaluate novel radar measurement synthesis with Peak Signal to Noise Ratio (PSNR). We additionally include the Structural Similarity Index (SSIM) to assess perceptual and structural fidelity, which complements PSNR by being more sensitive to local contrast and structural patterns, including the shapes and boundaries of dynamic targets. For occupancy estimation, we report Chamfer Distance (CD) and Relative Chamfer Distance (RCD) between predicted 2D bird’s eye view point clouds and LiDAR derived references. Since accurate ground-truth occupancy is difficult to extract directly from radar due to low azimuth resolution, multipath interference, and inherent sensor noise, we leverage synchronized LiDAR point clouds to construct geometric occupancy references. For each radar frame, we calibrate the corresponding LiDAR point clouds to radar coordinates using the provided LiDAR-to-radar transformation matrix. We then select points within the radar’s sensing range bounds and with 1<z<1-1<z<1. The resulting points are projected onto the $xy$-plane to generate the BEV ground-truth occupancy for that frame. To ensure compatibility with dynamic scenes, we perform occupancy evaluation using only the LiDAR point clouds corresponding to each radar timestamp, without accumulating across time.
@@ 5.2 Comparison with state-of-the-art
To ensure fair and rigorous evaluation, we focus exclusively on NF-based approaches, excluding methods relying on handcrafted environment modeling or sensor-specific signal simulations. We compare RF4D with three representative state-of-the-art methods: Radar Fields [4], D-NeRF [43], and Hexplane [7]. For Radar Fields, we directly adopt their official implementation, ensuring consistency and fair comparison. For D-NeRF and Hexplane, originally designed for RGB-based dynamic NVS, we adapt their architectures to our radar-based pipeline and retrain accordingly.
@@
Tables 1 and 2 present quantitative comparisons on the Oxford Radar RobotCar and Boreas datasets, respectively. RF4D consistently achieves the best performance across most driving scenarios on both datasets, outperforming all competing methods in both radar measurement synthesis (PSNR, SSIM) and occupancy estimation (CD, RCD). The jointly superior PSNR and SSIM results show that RF4D achieves both accurate and structurally consistent radar synthesis. Meanwhile, the lower CD and RCD values show that the predicted occupancies are spatially more precise and better aligned with LiDAR-derived geometry. In contrast, methods like D-NeRF and HexPlane, designed for RGB data without radar-specific modeling, yield inferior radar synthesis and occupancy estimation. Although Radar Fields incorporates radar sensing physics, its Bayesian filtering occupancy supervision is prone to multipath-induced false positives, which limit overall performance.
Figure 5: Radar robustness and generalization across weather conditions. While LiDAR point clouds degrade severely in snow, radar measurements remain stable. RF4D accurately reconstructs radar measurements, maintaining consistent performance across different weather conditions.
@@
Figure 4 presents qualitative comparisons of reconstructed radar measurements and occupancy estimation results across two representative scenes using Radar Fields [4] and our method. Our model successfully recovers dynamic objects (highlighted in red boxes), demonstrating its ability to capture temporal scene variations. We attribute this improvement to the incorporation of temporal information and temporal regularization, which enable RF4D to maintain consistent estimation of dynamic targets across frames. In addition, RF4D renders significantly cleaner radar reconstructions and predicts more accurate occupancy estimations that closely align with the LiDAR-derived ground truth. Radar measurements are often affected by sensor noise and multipath interference, which introduce false-positive responses. Radar Fields tends to misinterpret these spurious reflections as occupied regions, leading to false and unstable occupancy predictions. In contrast, RF4D employs a Gumbel–Sigmoid activation coupled with an occupancy gating mechanism during radar-specific power rendering, driving the neural field toward confident, near-binary occupancy outputs and effectively suppressing uncertain responses. Combined with the global occupancy sparsity regularization, these designs yield cleaner rendered radar measurements and more reliable occupancy maps. Figure 5 further verifies the radar robustness and generalization across weather conditions. While LiDAR point clouds become noisy and incomplete in snow conditions, radar measurements remain stable and reliable. Our RF4D framework accurately reproduces these radar measurements, showing consistent performance across different weather conditions.
@@ 5.3 Ablation study
In Table 3 and 4, we conduct ablation studies to analyze the contribution of each loss component. When trained with only the radar power reconstruction loss $\mathcal{L}_{rt}$, the model achieves acceptable PSNR but performs poorly in SSIM and occupancy metrics (CD and RCD). This is because, without occupancy constraints, the network tends to predict all regions as occupied ($\alpha\approx 1$) and overfit the radar power responses everywhere, resulting in inaccurate and structureless occupancy predictions. Introducing the occupancy regularization loss $\mathcal{L}_{p}$ encourages the model to focus on confident and stable regions, effectively suppressing uncertain occupancy estimates and improving both measurement synthesis and geometric accuracy. Note that when removing the temporal regularization losses $\mathcal{L}_{oc}$ and $\mathcal{L}_{m}$, we also remove the scene flow module $f_{\Delta x}$, since these components are designed to work together. Adding the temporal consistency loss $\mathcal{L}_{oc}$ further enforces temporal alignment of occupancy across frames by warping each point to its adjacent frames using the predicted motion offsets. However, this alone does not yield direct improvement, as the motion offsets can become unstable without explicit constraints. After further introducing the motion regularization term $\mathcal{L}_{m}$ to constrain the magnitude and smoothness of these offsets, the temporal predictions become more coherent, leading to consistent gains across all metrics.
5.4 Occupancy estimation comparison
In this section, we further compare our method with traditional radar occupancy estimation approaches, including Constant False Alarm Rate (CFAR) [45] and Bayesian filtering [62]. As shown in Table 5, our method consistently outperforms other approaches across most scenes. Conventional methods such as CFAR and Bayesian filtering often preserve multipath and noise artifacts, leading to unstable and inaccurate occupancy estimation. In contrast, our method provides accurate and stable occupancy estimates. This indicates that using occupancy maps from traditional methods as supervision during training may introduce misleading signals and hinder learning of precise scene structure. Notably, our approach does not rely on any external occupancy estimator for supervision, but learns directly from radar measurements.
Figure 6: 3D voxel grid reconstruction from radar measurements. Our method reconstructs full 3D occupancy geometry from sparse and low-resolution radar data, capturing both moving vehicles and static objects present in the scene.
5.5 3D voxel grid reconstruction
We adopt physics-based importance sampling [4], which uses radar antenna gain profiles to obtain elevation information to further reconstruct 3D geometry. As shown in Figure 6, our RF4D with physics-based importance sampling can reconstruct full 3D occupancy geometry of the scene from sparse and low-resolution radar data, successfully capturing both moving vehicles and static objects in the scene. However, it is important to note that the elevation field of view (FoV) of the radar is limited to around 2 degrees, which restricts the reconstruction ability for objects located outside the radar’s elevation coverage.